Based on a​ poll, 69​% of Internet users are more careful about personal information when using a public​ Wi-Fi hotspot. What is the probability that among four randomly selected Internet​ users, at least one is more careful about personal information when using a public​ Wi-Fi hotspot? How is the result affected by the additional information that the survey subjects volunteered to​ respond?



1
Expert's answer
2021-06-07T01:41:53-0400

This is a binomial distribution with n=4, p=0.69.

"P(X\ge1)=1-P(X=0)=1-(1-0.69)^4=0.9908."

Volunteered respond produces a bias. Thus, above result would not be relevant.